OpenStudy (anonymous):

Give the equation of the line passing through the point (1,-5) that is parallel to y = 12 - 8x ???

OpenStudy (campbell_st):

whats the slope of the given line..?

OpenStudy (akashdeepdeb):

If a line is parallel to another line..Then both have the same slope!! :D Can you find the slope of this line? By using the formula y=mx+b [Where m is the slope] :D

OpenStudy (akashdeepdeb):

@Find_Jeff ?

OpenStudy (anonymous):

I don't understand ? :c

OpenStudy (campbell_st):

lol.... well we don't have a multiple choice list... the given line is y = -8x + 12 the slope intercept form is y = mx + b m is the slope so by comparing what is the value of m...?

OpenStudy (anonymous):

-8, because it would be the same, right?

OpenStudy (campbell_st):

thats correct so the line will be y = -8x + b you know the point.... x = 1 and y = -5 substitute them yo find the value of b, which is the y-intercept

OpenStudy (anonymous):

OH alright! I get it! :D thank you so much!

OpenStudy (campbell_st):

glad to help...